# In Search of Lost Online Test-time Adaptation: A Survey

### v2

@@ Abstract
This paper presents a comprehensive survey on online test-time adaptation (OTTA), a paradigm focused on adapting machine learning models to novel data distributions upon batch arrival. Despite the recent proliferation of OTTA methods, the field is mired in issues like ambiguous settings, antiquated backbones, and inconsistent hyperparameter tuning, obfuscating the real challenges and making reproducibility elusive. For clarity and a rigorous comparison, we classify OTTA techniques into three primary categories and subject them to benchmarks using the potent Vision Transformer (ViT) backbone to discover genuinely effective strategies. Our benchmarks span conventional corrupted datasets such as CIFAR-10///100-C and ImageNet-C and real-world shifts embodied in CIFAR-10.1 and CIFAR-10-Warehouse, encapsulating variations across search engines and synthesized data by diffusion models. To gauge efficiency in online scenarios, we introduce novel evaluation metrics, including GFLOPs, shedding light on the trade-offs between adaptation accuracy and computational overhead. Our findings diverge from existing literature, indicating: (1) transformers exhibit heightened resilience to diverse domain shifts, (2) the efficacy of many OTTA methods hinges on ample batch sizes, and (3) stability in optimization and resistance to perturbations are critical during adaptation, especially when the batch size is 1. Motivated by these insights, we pointed out promising directions for future research.
@@ 1 Introduction
Dataset shift (Quinonero-Candela et al, 2008) poses a notable challenge for machine learning. Models often experience significant performance drops when confronting test data characterized by conspicuous distribution differences from training. Such differences might come from changes in style and lighting conditions and various forms of corruption, making test data deviate from the data upon which these models were initially trained. To mitigate the performance degradation during inference, test-time adaptation (TTA) has emerged as a promising solution. TTA aims to rectify the dataset shift issue by adapting the model to novel distributions using unlabeled test data (Liang et al, 2023). Different from the traditional paradigm of domain adaptation (Ganin and Lempitsky, 2015; Wang and Deng, 2018), TTA does not require access to source data for distribution alignment. Commonly used strategies in TTA include unsupervised proxy objectives, spanning techniques such as pseudo-labeling Liang et al (2020), graph-based learning (Luo et al, 2023), and contrastive learning Chen et al (2022), applied on the test data through multiple training epochs to enhance model accuracy, such as autonomous vehicle detection (Hegde et al, 2021), pose estimation (Lee et al, 2023), video depth prediction (Liu et al, 2023a), frame interpolation (Choi et al, 2021), and medical diagnosis (Ma et al, 2022; Wang et al, 2022b; Saltori et al, 2022). Nevertheless, requiring access to the complete test set at every time step may not always align with practical use. In many applications, such as autonomous driving, adaptation is restricted to using only the current test batch processed in a streaming manner. Such operational restrictions make it untenable for TTA to require the full test set tenable.
In this study, our focus is on a specific line of TTA methods, i.e., online test-time adaptation (OTTA), which aims to accommodate real-time changes in the test data distribution. We provide a comprehensive overview of existing OTTA studies and evaluate the efficiency and effectiveness of these methods and their individual components. To facilitate a structured comprehension of the OTTA landscape, we categorize existing approaches into three groups: data-centric OTTA, model-based OTTA, and optimization-oriented OTTA.
@@
  • Model-based OTTA changes the original backbone, such as modifying specific layers or their mechanisms, adding supplementary branches, and incorporating prompts.
  • Optimization-based OTTA focuses on various optimization methods. Examples are designing new loss functions, updating BatchNorm layers during testing, using pseudo-labeling strategies, teacher-student frameworks, and contrastive learning-based approaches, etc.
It is potentially useful to combine methods from different categories for further improvement. An in-depth analysis of this strategy is presented in Sec. 3. Note that this survey does not include the paper if source-stage customization is needed, such as (Thopalli et al, 2022; Döbler et al, 2023; Brahma and Rai, 2023; Jung et al, 2022; Adachi et al, 2023; Lim et al, 2023; Choi et al, 2022; Chakrabarty et al, 2023; Marsden et al, 2022; Gao et al, 2023)
Differences from an existing survey. Liang et al (2023) provides a comprehensive overview of the vast topic of test-time adaptation (TTA), discussing TTAs in diverse configurations and their applicability in vision, natural language processing (NLP), and graph data analysis. One limitation is that the survey does not provide experimental comparisons of existing methods. Mounsaveng et al (2023) studies fully test-time adaptation for some specific items, e.g., batch normalization, calibration, class re-balancing, etc. However, it does not focus too much on analyzing the existing methods and exploring ViTs. In contrast, our survey concentrates on online TTA approaches and provides valuable insights from experimental comparisons, considering hyperparameter selection and backbone influence (Zhao et al, 2023b). Contributions. In particular, with the ascent of vision transformer (ViT) architectures in various machine learning domains, this survey studies whether OTTA strategies developed for ResNet structures maintain their effectiveness after being integrated into ViT instead. To this end, we benchmark seven state-of-the-art OTTA algorithms on a wide range of distribution shifts under a new set of evaluation metrics. Below, we summarize the key contribution of this survey.
@@
  • [Benchmarking OTTA strategies with ViT] We reimplemented representative OTTA baselines under the ViT architecture and testified their performance against five benchmark datasets. We drive a set of replacement rules that adapt the existing OTTA methods to accommodate the new backbone.
  • [Both accuracy and efficiency as evaluation Metrics] Apart from using the traditional recognition accuracy metric, we further provide insights into various facets of computational efficiency by Giga floating-point operations per second (GFLOPs). These metrics are important in real-time streaming applications.
  • [Real-world testbeds] While existing literature extensively explores OTTA methods on corruption datasets like CIFAR-10-C, CIFAR-100-C, and ImageNet-C, our interests are more fall into their capability to navigate real-world dataset shifts. Specifically, we assess OTTA performance on CIFAR-10-Warehouse, a newly introduced, expansive test set of CIFAR-10. Our empirical analysis and evaluation lead to different conclusions from findings in the existing survey (Liang et al, 2023).
This work aims to summarize existing OTTA methods with the aforementioned three categorization criteria and analyze some representative approaches using empirical results. Moreover, to assess real-world potential, we conduct comparative experiments to explore the portability, robustness, and environmental sensitivity of the OTTA components. We expect this survey to offer a systematic perspective in navigating OTTA’s intricate and diverse solutions, enabling a clear identification of effective components. We also present new challenges as potential future research directions.
@@
Datasets ##\## domains ##\## test images ##\## classes corrupted? image size
@@
ImageNet-C (Hendrycks and Dietterich, 2018) 19 4750,000 1000 Yes $224\times 224$
@@
Figure 1: Taxonomy of existing OTTA methods. The categories, i.e., optimization-based, data-based, and model-based, inform three mainstream working mechanisms. Prompt-related methods are categorized into model-based methods.
Organization of the survey. The rest of this survey will be organized as follows. Sec. 2 presents the problem definition and introduces widely used datasets, metrics, and applications. Using the taxonomy shown in Fig. 1, Sec. 3 comprehensively reviews existing OTTA methods. Then, using transformer backbones, Sec. 4 empirically analyzes seven state-of-the-art methods based on new evaluation metrics on corrupted and real-world distribution shifts. We conclude the survey in Sec. 6.
@@ 2.1 Problem Definition
In OTTA, we assume access to a trained source model and adapt the model at test time over the test input before making the final prediction. The given source model $f_{\theta^{S}}$ parameterized by $\theta^{S}$ is pre-trained on a labeled source domain $\mathcal{D}_{S}=\{(\bm{x}^{S},\bm{y}^{S})\}$, which is formed by i.i.d. sampling from the source distribution $p_{S}$. Unlabeled test data $\mathcal{D}_{T}=\{\bm{x}_{1}^{T},\bm{x}_{2}^{T},\bm{x}_{t}^{T},\ldots,\bm{x}_{n}^{T}\}$ come in batches, where $t$ indicates the current time step, $n$ is the overall time steps (i.e., number of batches). Test data often come from one or multiple different distributions $(\bm{x}_{t}^{T},\bm{y}_{t}^{T})\sim p_{T}$, where $p_{S}(\bm{x},y)\neq p_{T}(\bm{x},y)$ under the covariate shift assumption (Huang et al, 2006). During TTA, we update the model parameters batch $t$, resulting in an adapted model $f_{\theta^{t}}$. Before adaptation, the pre-trained model is expected to retain its original architecture, including the backbone, without modifying its layers or introducing new model branches during training. Additionally, the model is restricted to observing the test data only once and must produce predictions promptly. By refining the definition of OTTA in this manner, we aim to minimize limitations associated with its application in real-world settings. Note that following adaptation to a specific domain, the model is reset to its original pretrained state. i.e., $\underline{f_{\theta^{S}}}\rightarrow f_{\theta^{0}}\rightarrow\underline{f_{\theta^{S}}}\rightarrow f_{\theta^{1}}\rightarrow\underline{f_{\theta^{S}}}\rightarrow\cdots f_{\theta^{t}}$.
Due to the covariate shift between the source and test data, adapting the source model without the source data poses a significant challenge. Since there is no way to align these two sets, one may ask, what kind of optimization objective could work under such a limited environment? Meanwhile, as the test data come at a fixed pace, how many could be a desirable amount to best fix the test-time adaptation? Or will the adaptation even work in the new era of backbones (e.g., ViTs)? Does “Test-time Adaptation” become a false proposition with the backbone upgrading? With these concerns, we unfold the OTTA methods by their datasets, evaluations, and applications and decouple their strategies, aiming to discover which one and why one component would work with the update of existing backbones.
@@ 2.2 Datasets
This survey mainly summarizes datasets in image classification, a fundamental problem in computer vision, while recognizing that OTTA has been applied to many downstream tasks (Ma et al, 2022; Ding et al, 2023; Saltori et al, 2022). Testbeds in OTTA usually seek to facilitate adaptation from natural images to corrupted ones. The latter are created by perturbations such as Gaussian noise and Defocus blur. Despite including corruptions at varying severities, these synthetically induced corruptions may not sufficiently mirror the authentic domain shift encountered in real-world scenarios. Our work uses corruption and real-world shift datasets, summarized in Table 1. Details of each testbed are described below.
  • CIFAR-10-C is a standard benchmark for image classification. It contains 950,000 color images, each of 32x32 pixels, spanning ten distinct classes. CIFAR-10-C retains the class structure of CIFAR-10 but incorporates 15 diverse corruption styles, with severities ranging from levels 1 to 5. This corrupted variant aims to simulate realistic image distortions or corruptions that might arise during processes like image acquisition, storage, or transmission.
@@
  • ImageNet-C is a corrupted version of ImageNet test set (Krizhevsky et al, 2012). Produced from ImageNet-1k, ImageNet-C has $19$ types of corruption domains, including $4$ validation corruptions. For each domain, $5$ levels of severity are produced, with $50,000$ images per severity from $1,000$ classes.
Although testbeds with corruptions have been extensively utilized, they might not fully capture the complexities of real-world scenarios as they represent artificially created domain differences. Experimental benchmarking from the real world is still lacking. Therefore, this paper evaluates OTTA on real-world test sets such as CIFAR-10-Warehouse and CIFAR-10.1 to address the limitations of such testbeds.
  • CIFAR-10.1 (Recht et al, 2018) is a real-world test set with the same label space as CIFAR-10. It contains roughly 2,000 images sampled from the Tiny Image dataset (Yang et al, 2016).
  • CIFAR-10-Warehouse (CIFAR-10-W) integrates images from both diffusion models, specifically Stable-diffusion-2-1 (Rombach et al, 2022), and targeted keyword searches across seven popular search engines. Comprising $37$ generated and $143$ real-world datasets, each subset has $300$ to $8,000$ images, revealing noticeable within-class variations across different search criteria.
@@ 2.3 Evaluation
Number of updated parameters provides insights into the complexity of the adaptation process. A model that requires a large number of updated parameters may not be practical for online adaptation.
@@ 2.4 Relationship with Other Tasks
Offline test-time adaptation (TTA), also called source-free domain adaptation (Liang et al, 2020, 2022; Ding et al, 2022; Yang et al, 2021), is a technique to adapt a source pre-trained model to the target (i.e., test) set. This task assumes that the model can access the entire dataset multiple times. This differs from online test-time adaptation, where the test data is given in batches.
Continual TTA While OTTA requires resetting the adapted model back to the source pre-trained one for every distinct domain (i.e., corruption type), continual TTA (e.g., (Wang et al, 2022a; Hong et al, 2023; Song et al, 2023; Chakrabarty et al, 2023; Gan et al, 2023)) assumes no domain information available and does not allow any reset mechanism. Although the corruption domains in continuous TTA appear one by one, with clear boundaries between each domain, there is no separate indication of domain information during adaptation.
Gradual TTA tackles real-world scenarios where domain shifts are gradually introduced through incoming test samples (Marsden et al, 2022; Döbler et al, 2023). An example is the gradual and continuous change in weather conditions. For corruption datasets, existing gradual TTA approaches assume that test data transition from severity level 1 to level 2 and then progress slowly toward the highest level. Continual and gradual TTA methods also support online test-time adaptation (i.e., episodic learning).

### v3

# Corresponding author

@@ Abstract
This article presents a comprehensive survey of online test-time adaptation (OTTA), focusing on effectively adapting machine learning models to distributionally different target data upon batch arrival. Despite the recent proliferation of OTTA methods, conclusions from previous studies are inconsistent due to ambiguous settings, outdated backbones, and inconsistent hyperparameter tuning, which obscure core challenges and hinder reproducibility. To enhance clarity and enable rigorous comparison, we classify OTTA techniques into three primary categories and benchmark them using a modern backbone, the Vision Transformer (ViT). Our benchmarks cover conventional corrupted datasets such as CIFAR-10/100-C and ImageNet-C, as well as real-world shifts represented by CIFAR-10.1, OfficeHome, and CIFAR-10-Warehouse. The CIFAR-10-Warehouse dataset includes a variety of variations from different search engines and synthesized data generated through diffusion models. To measure efficiency in online scenarios, we introduce novel evaluation metrics, including GFLOPs, wall clock time, and GPU memory usage, providing a clearer picture of the trade-offs between adaptation accuracy and computational overhead. Our findings diverge from existing literature, revealing that (1) transformers demonstrate heightened resilience to diverse domain shifts, (2) the efficacy of many OTTA methods relies on large batch sizes, and (3) stability in optimization and resistance to perturbations are crucial during adaptation, particularly when the batch size is 1. Based on these insights, we highlight promising directions for future research. Our benchmarking toolkit and source code are available at https://github.com/Jo-wang/OTTA_ViT_survey.
@@ 1 Introduction
Dataset shift (Quinonero-Candela et al, 2008) poses a notable challenge for machine learning. Models often experience significant performance drops when confronting test data characterized by superior distribution differences from training. Such differences might come from changes in style and lighting conditions and various forms of corruption, making test data deviate from the data upon which these models were initially trained. To mitigate the performance degradation brought by these shifts, test-time adaptation (TTA) has emerged as a promising solution. TTA aims to rectify the dataset shift issue by adapting the model to unseen distributions using unlabeled test data (Liang et al, 2023). Different from unsupervised domain adaptation (Ganin and Lempitsky, 2015; Wang and Deng, 2018), TTA does not require access to source data for distribution alignment. Commonly used strategies in TTA include unsupervised proxy objectives, spanning techniques such as pseudo-labeling Liang et al (2020), graph-based learning (Luo et al, 2023), and contrastive learning Chen et al (2022), applied on the test data through multiple training epochs to enhance model accuracy, such as autonomous vehicle detection (Hegde et al, 2021), pose estimation (Lee et al, 2023), video depth prediction (Liu et al, 2023a), frame interpolation (Choi et al, 2021), and medical diagnosis (Ma et al, 2022; Wang et al, 2022b; Saltori et al, 2022). Nevertheless, requiring access to the complete test set at every time step may not always align with practical use. In many applications, such as autonomous driving, adaptation is restricted to using only the current test batch processed in a streaming manner. Such operational restrictions make it untenable for TTA to require the full test set tenable.
In this study, our focus is on a specific line of TTA methods, i.e., online test-time adaptation (OTTA), which aims to accommodate real-time changes in the test data distribution. We provide a comprehensive overview of existing OTTA studies and evaluate the efficiency and effectiveness of these methods and their individual components. To facilitate a structured comprehension of the OTTA landscape, we categorize existing approaches into three groups: optimization-oriented OTTA, data-centric OTTA, and model-based OTTA.
  • Optimization-based OTTA focuses on various optimization methods. Examples are designing new loss functions, updating normalization layers (e.g., BatchNorm) during testing, using pseudo-labeling strategies, teacher-student frameworks, and contrastive learning-based approaches, etc.
@@
  • Model-based OTTA adjusts the original backbone, such as modifying specific layers or their mechanisms, adding supplementary branches, and incorporating prompts.
It is potentially useful to combine methods from different categories for further improvement. An in-depth analysis of this strategy is presented in Sec. 3. Note that this survey does not include the paper if source-stage customization is needed, such as Thopalli et al (2022); Döbler et al (2023); Brahma and Rai (2023); Jung et al (2022); Adachi et al (2023); Lim et al (2023); Choi et al (2022); Chakrabarty et al (2023); Marsden et al (2022); Gao et al (2023). As it necessitates extra operation or information at the source pre-training stage, this setup may not be feasible in some scenarios and may prevent us from comparing on a consistent standard.
Differences from an existing survey. Liang et al (2023) provided a comprehensive overview of the vast topic of test-time adaptation (TTA), discussing TTAs in diverse configurations and their applicability in vision, natural language processing (NLP), and graph data analysis. One limitation is that the survey did not provide experimental comparisons of existing methods. Mounsaveng et al (2023) studied fully test-time adaptation for some specific components, e.g., batch normalization, calibration, class re-balancing, etc. However, it did not focus too much on analyzing the existing methods and exploring ViTs. Marsden et al (2024) conducted comprehensive study on universal test-time adaptation setting. In contrast, our survey concentrates on purely online TTA approaches and provides valuable insight from experimental comparisons, considering various domain shifts, hyperparameter selection, and backbone influence (Zhao et al, 2023b).
Contributions. As Vision Transformer (ViT) architectures gain increasing prominence, a critical question emerges: As Vision Transformer (ViT) architectures gain increasing prominence, a critical question arises: Do OTTA strategies, originally devised for CNNs, retain their effectiveness when applied to ViT models? This question stems from the significant architectural differences between ViTs and conventional CNNs, such as ResNets, particularly in their normalization layers and information processing mechanisms. Given the growing adoption of ViTs, investigating their compatibility with existing OTTA strategies is essential. To thoroughly explore this question, we evaluate eight representative OTTA algorithms across diverse distribution shifts, employing a set of metrics to evaluate both effectiveness and efficiency. Below, we summarize the key contribution of this survey:
@@
  • [Comprehensive Benchmarking and Adaptation of OTTA Strategies with ViT] We reimplemented representative OTTA baselines under the ViT architecture and testified their performance against six benchmark datasets. We drive a set of replacement rules that adapt the existing OTTA methods to accommodate the new backbone.
  • [Both accuracy and efficiency as evaluation Metrics] Apart from using the traditional recognition accuracy metric, we further provide insights into various facets of computational efficiency by Giga floating-point operations per second (GFLOPs), wall clock time, and GPU memory usage. These metrics are important in real-time streaming applications and can be treated as a supplementary of (Marsden et al, 2024).
  • [Real-world testbeds] While existing literature extensively explores OTTA methods on corruption datasets like CIFAR-10-C, CIFAR-100-C, and ImageNet-C, or even real-world datasets (Marsden et al, 2024) our interests are more fall into their capability to navigate real-world dataset shifts from different angles. Specifically, we assess OTTA performance on CIFAR-10-Warehouse, a newly introduced, expansive test set of CIFAR-10. Conclusions about various domain shifts, based on the same pre-trained model, are drawn from our empirical analysis and evaluation. These shifts were previously unexplored in the existing survey (Liang et al, 2023).
This work aims to summarize existing OTTA methods with the aforementioned three categorization criteria and analyze some representative approaches by empirical results. Moreover, to assess real-world potential, we conduct comparative experiments to explore the portability, robustness, and environmental sensitivity of the OTTA components. We expect this survey to offer a systematic perspective in navigating OTTA’s intricate and diverse solutions, enabling a clear identification of effective components. We also present new challenges as potential future research directions.
@@
Datasets ##\## domains ##\## images ##\## classes corrupted? image size
@@
ImageNet-C (Hendrycks and Dietterich, 2018) 19 4,750,000 1000 Yes $224\times 224$
@@
OfficeHome (Venkateswara et al, 2017) 4 15,500 65 No non-uniform
Figure 1: Taxonomy of existing OTTA methods. The categories, i.e., optimization-based, data-based, and model-based, inform three mainstream working mechanisms. To provide a clear illustration, methods related to prompts are categorized into model-based methods.
Organization of the survey. The rest of this survey will be organized as follows. Sec. 2 presents the problem definition and introduces widely used datasets, metrics, and applications. Using the taxonomy shown in Fig. 1, Sec. 3 comprehensively reviews existing OTTA methods. With the new generation of backbones, Sec. 4 empirically analyzes eight state-of-the-art methods based on multiple evaluation metrics on corrupted and real-world distribution shifts. We introduce the potential future directions in Sec. 5 and conclude this survey in Sec. 6.
@@ 2.1 Problem Definition
In OTTA, we assume access to a trained source model and adapt the model at test time over the test input to make the final prediction. The given source model $f_{\theta^{S}}$ parameterized by $\theta^{S}$ is pre-trained on a labeled source domain $\mathcal{D}_{S}=\{(\bm{x}^{S},\bm{y}^{S})\}$, which is formed by i.i.d. sampling from the source distribution $p_{S}$. Unlabeled test data $\mathcal{D}_{T}=\{\bm{x}_{1}^{T},\bm{x}_{2}^{T},\bm{x}_{t}^{T},\ldots,\bm{x}_{n}^{T}\}$ come in batches, where $t$ indicates the current time step, $n$ is the total number of time steps (i.e., number of batches). Test data often come from one or multiple different distributions $(\bm{x}_{t}^{T},\bm{y}_{t}^{T})\sim p_{T}$, where $p_{S}(\bm{x},y)\neq p_{T}(\bm{x},y)$ under the covariate shift assumption (Huang et al, 2006). During TTA, we update the model parameters for batch $t$, resulting in an adapted model $f_{\theta^{t}}$. Before adaptation, the pre-trained model is expected to retain its original architecture, especially the backbone, without modifying its layers or introducing new model branches during training. Additionally, the model is restricted to observing the test data only once and must produce predictions promptly online. By refining the definition of OTTA in this manner, we aim to minimize limitations associated with its application in real-world settings. Note that the model is reset to its original pre-trained state after being adapted to a specific domain. i.e., $\underline{f_{\theta^{S}}}\rightarrow f_{\theta^{0}}\rightarrow\underline{f_{\theta^{S}}}\rightarrow f_{\theta^{1}}\rightarrow\underline{f_{\theta^{S}}}\rightarrow\cdots f_{\theta^{t}}$.
Due to the covariate shift between the source and test data, adapting the source model without the source data poses a significant challenge. Since there is no way to align these two sets as in unsupervised domain adaptation, one may ask, what kind of optimization objective could work under such a limited environment? Meanwhile, as the test data come at a fixed pace, how many could be a desirable amount to best fix the test-time adaptation? Or will the adaptation even work in the new era of backbones (e.g., ViTs)? Does “Test-time Adaptation” become a false proposition with the backbone changes? With these concerns, we unfold the OTTA methods by their datasets, evaluations, and applications and decouple their strategies, aiming to discover which one and why one component would work with the update of existing backbones.
@@ 2.2 Datasets
This survey mainly summarizes datasets in image classification, a fundamental problem in computer vision, while recognizing that OTTA has been applied to many downstream tasks (Ma et al, 2022; Ding et al, 2023; Saltori et al, 2022). Testbeds in OTTA usually seek to facilitate adaptation from natural images to corrupted ones. The latter are created by perturbations such as Gaussian noise and Defocus blur. Despite including corruptions at varying severities, these synthetically induced corruptions may not sufficiently mirror the authentic domain shift encountered in real-world scenarios. Our work uses corruption (Croce et al, 2021), generated images, and real-world shift datasets, summarized in Table 1. Details of each testbed are described below.
CIFAR-10-C is a standard benchmark for image classification. It contains 950,000 color images, each of 32x32 pixels, spanning ten distinct classes. CIFAR-10-C retains the class structure of CIFAR-10 but incorporates 19 diverse corruption styles, with severities ranging from levels 1 to 5. This corrupted variant aims to simulate realistic image distortions or corruptions that might arise during processes like image acquisition, storage, or transmission.
@@
ImageNet-C is a corrupted version of ImageNet test set (Krizhevsky et al, 2012). Produced from ImageNet-1k, ImageNet-C has a similar setup to the CIFAR-10-C and CIFAR-100-C corruption types. For each domain, $5$ levels of severity are produced, with $50,000$ images per severity from $1,000$ classes.
CIFAR-10.1 (Recht et al, 2018) is a real-world test set of CIFAR-10. It contains roughly 2,000 images sampled from the Tiny Image dataset (Yang et al, 2016).
CIFAR-10-Warehouse (Sun et al, 2023) integrates images from both diffusion model (i.e., Stable Diffusion-2-1 (Rombach et al, 2022)) and targeted keyword searches across eight popular search engines. The diffusion model uses the prompt “high quality photo of color class name”, with color chosen from 12 options. The dataset comprises 37 generated and 143 real-world subsets, each containing between 300 and 8,000 images.
OfficeHome is a widely used benchmark in domain adaptation and domain generalization tasks. It has $65$ classes within $4$ domains: Artistic images (Art), Clip Art, Product images (Product), and Real-World images (RealWorld). Each domain represents a significantly different style of images.
@@ 2.3 Evaluation
Wall-clock time measures the actual time taken by the model to complete the adaptation. This may significantly influence the use case of an OTTA model.
GPU memory usage refers to the amount of memory the model uses while running on a GPU. A model with lower GPU memory usage is more applicable and can be deployed on a wider range of devices.
@@ 2.4 Relationship with Other Tasks
Offline test-time adaptation (TTA) (Liang et al, 2020, 2022; Ding et al, 2022; Yang et al, 2021) is a technique to adapt a source pre-trained model to the target (i.e., test) set. This task assumes that the model has access to the entire dataset multiple times, which differs from the online test-time adaptation where the test data is given in batches.
Continual TTA Contrary to the classic OTTA setup, where adaptation could be envisioned as occurring in discrete steps corresponding to distinct domain shifts, continual TTA (Wang et al, 2022a; Hong et al, 2023; Song et al, 2023; Chakrabarty et al, 2023; Gan et al, 2023) operates under the premise of seamless, continuously adapting to new data distributions. This process does not need a reset of the model with each perceived domain shift. Rather, it highlights the importance of a model’s ability to autonomously update and refine its parameters in response to ongoing changes in the data landscape, without explicit indicators of domain boundaries.
Gradual TTA tackles real-world scenarios where domain shifts are gradually introduced through incoming test samples (Marsden et al, 2022; Döbler et al, 2023). An example is the gradual and continuous change in weather conditions. For corruption datasets, existing gradual TTA approaches assume that test data transition from severity level 1 to level 2 and then progress slowly toward the highest level. Continual and gradual TTA methods also support online test-time adaptation.